OpenStudy (anonymous):

MEDAL FOR RIGHT ANSWER!!! PLEASE HELP!!! Thomas walks his dog for 15 minutes today-5 more minutes than yesterday. Express as a percent his walking time today compared to his walking time yesterday. round to the nearest percent.

OpenStudy (anonymous):

@Abhisar

OpenStudy (anonymous):

@lexie<3

OpenStudy (anonymous):

I don't know but someone else had the same question so this link might help http://widget.openstudy.com/study#/updates/5494481ce4b01babb0df77d6

OpenStudy (mathmate):

Hint: Example: I ate 20 cookies yesterday. Today I ate 30. What is the percentage of cookies I ate today compared to what I ate yesterday? First we find the base value which is considered 100%. Question says "compared to what I ate yesterday", so what I ate yesterday (20) is the base percent of 100%. After that, we form a proportion in fraction form, let x=percentage I ate today "compared to yesterday". \( \dfrac{\#cookies~today}{\#cookies~yesterday}=\dfrac{x}{100\%} \) Now you can solve for x after substituting the number of cookies. \( \dfrac{30}{20}=\dfrac{x}{100\%} \) Cross multiply to get \(x=\dfrac{30}{20}*100\% =150\%\)